OpenStudy (anonymous):

does anybody know how to solve this? 3|-4x+75|+9=144 i want the steps to doing it please

OpenStudy (anonymous):

3|-4x+75|+9=144 3|-4x+75|=135 |-4x+75|=45 -4x + 75 = 45 4x - 75 = 45 x = 30 / 4 = 15/2 x = 120 / 4 = 30 \[x = \frac{15}{2}, 30\]